OpenStudy (anonymous):

[(5^6n x 9^4n x 15^2n) / 3^2n] = k^8n Find the value of k where k is a positive integer. Show me the steps thanks, no straight answers

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

err is 5^6n not 56^6n

OpenStudy (anonymous):

\[\frac{5^{6n}∗9^{4n}∗15^{2n}}{3^{2n}}=k^{8n}\]

OpenStudy (anonymous):

See, 9 = 3*3 = 3^2 15 = 5*3 \[\frac{5^{6n}*3^{8n}*3^{2n}*5^{2n}}{3^{2n}} = 5^{6n +2n}*3^{8n} = 5^{8n}*3^{8n}\] \[(5*3)^{8n} = k^{8n}\] So, can you find k from it or not???

OpenStudy (anonymous):

ok thanks alot <3

OpenStudy (anonymous):

Welcome dear...